OpenStudy (anonymous):

inverse f(x) = ln(x/x-1)

OpenStudy (anonymous):

Maybe this will help The inverse of the natural log is \[e^{x}\]

OpenStudy (anonymous):

y=ln(x/x-1), e^y=x/x-1 (e^y)x- e^y=x x= e^y/(e^y-1). now plug x at the place of y. inverce of F(x)= (e^x)/(e^x-1)